OpenStudy (anonymous):

Need help: What is the value of y in the solution to the following system of equations? 2x + y = −4 5x + 3y = −6

OpenStudy (luigi0210):

Any specific method you have in mind?

OpenStudy (anonymous):

No idea. Lol.

OpenStudy (anonymous):

But I am guessing y is the same for each equation, because their both in the same question.

OpenStudy (luigi0210):

Yup, so the best way, if you ask me, is to use elimination.

OpenStudy (anonymous):

Elimination, as in use random inputs for y and x?

OpenStudy (luigi0210):

No, as in getting rid of one variable to get the other

OpenStudy (anonymous):

y = -2(x + 2)

OpenStudy (anonymous):

That's for 2x + y = - 4

OpenStudy (anonymous):

y = 1/3(-5x - 6) for 5x + 3y = -6

OpenStudy (luigi0210):

You don't need to factor them Mikeyy :P

OpenStudy (anonymous):

Well, now I have no idea again.

OpenStudy (luigi0210):

This is what I mean by Elimination: http://www.mathportal.org/algebra/solving-system-of-linear-equations/elimination-method.php

OpenStudy (anonymous):

Well, with 5x + 3y= -6 what would I do to eliminate?

OpenStudy (luigi0210):

The way the equation is set up is fine, now what we need to do is cancel x out so we have y left over. To do this we need to make x have a value that is the same.. but one being positive and the other being negative, like this one:  You see how the values canceled and we could solve for the other easily?